At a certain temperature, the Kp for the decompositon of H2S is 0.834.

H2S(g) <---> H2(g) + S(g)

Initally, only H2S is present at a pressure of 0.108 bar in a closed container. What is the total pressure in the container at equilibrium?
Answer #1

-0.834 + 0.695 +0.36 hs(g) → H₂(g) + sig t=0 0.108 eq bn 0.108-P P P kp = PH = -0.834 + V1.055 IS = -0.834 + 1.0271 2. 0193 =